OpenStudy (anonymous):

find the amount of interest earned on an investment: principal of $200, interest rate of 4 percent per year compounded quarterly, period of 2 yrs

hero (hero):

Oh, okay...compound interest... :)

hero (hero):

So, how would you like to be helped? Do you want someone to explain to you how to solve it or are you just here to check your answer?

OpenStudy (anonymous):

I checked the solution manual and I found 200 ((1.01) to the 8th power) - 1. My question is how did they get 1.01

hero (hero):

Well, okay, I can type out the solution for you...

hero (hero):

The formula for compound interest quarterly is: \[A = P(1+\frac{r}{n})^{nt} \] So plugging in the known values of P, n, and t, we get: \[A = 200(1+\frac{.04}{4})^{4*2}\] So now, just following simple order of operations...we deal with the parentheses first, and within the parentheses, division has priority, so if you divide .04/4 you get .01. Add 1 to that and you get 1.01: \[A = 200(1.01)^{8} \] Now, where you got the -1 from...you'll have to explain that one to me.

OpenStudy (anonymous):

Thank you you are a true angel...now could you also enlighten me by telling me how can you tell when to use the compound interest versus the basic formula?

hero (hero):

What basic formula?

hero (hero):

First of all....I notice one thing about compound interest problems...They almost ALWAYS say the words "compounded quarterly" or compounded annually" or "compounded daily" or compunded continuously"... as soon as I see the word "compounded" I think to myself, "Oh, this is a compound interest problem.". Problems that involve "basic formulas" probably wouldn't have "compounding interest".

hero (hero):

I hope that helps.